OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $78 for 1 adult ticket and 2 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 78 = one-halfx + (x - 6) What is the regular price of a child's ticket? $50 $56 $75 $81 @TheRealMeeeee @katalina13

OpenStudy (therealmeeeee):

Regularly: c = a-6 and a = a But this is during the off-peak season, so everything is half off: c = (a-6)/2 and a = a/2 So the equation for off-peak is: 78 = (a/2) + 2((a-6)/2) {the two 2s on the right side of the plus sign will cancel each other out} 78 = (a/2) + (a - 6) Remove parentheses through distributing and then add 6 to both sides: 84 = a/2 + a 84 = a(1/2 + 1) 84 = a(1.5) Divide both sides by 1.5 a = 56 which is the adult ticket price. Subtract 6 from that to get the child's ticket price: $50.

OpenStudy (anonymous):

got it thanks